2018 April

From EOVSA Wiki
Jump to navigation Jump to search

Back to Observing Log

Apr 02

14:01 UT This morning after the low-frequency receiver observation, I updated the Ant14-lo delay by 4 ns, so the solar observations will differ from the phase calibration by that amount, for the low-frequency receiver calibration.

19:20 UT Oops. I inadvertently updated the high-frequency receiver delay on Ant 14 this morning. This does not affect solar observations, but did mess up the phasecals for today. I have now updated both high- and low-receiver delays. My high-receiver changes affect more than just Ant 14, since many of the delays were off by 0.1 ns from their optimal value, so it will be simplest just to use the end-of-the-day refcal for calibration of today's data.

Apr 06

19:00 UT I received an email from Mark at OVRO saying that the Ant 14 receiver had warmed up, which turns out to be due to a vacuum leak in one of the windows (the Hi-freq Receiver, I guess, since the other window is not under vacuum). He will see if the window can be repaired on the telescope, or must be brought down to fix. Meanwhile, we will have no calibrations for the foreseeable future.

Apr 09

13:11 UT I discovered this morning that the UDB processing pipeline was not running since Apr. 06. I removed the lock file, and although it shows signs of working, no new files are created yet.

Apr 10

15:00 UT The 27-m antenna is back, although still cooling down. Both HPol and VPol on the low-frequency receiver appear to be dead, though.

Apr 15

21:16 UT The correlator was showing signs of bad sync again for the last couple of days. I rebooted the correlator at about 20:10 UT, and then measuring delays on 3C84. The network was glitching, though, so it took a little longer than I wanted. I just now got the delays reset, and back on the Sun. There was about 30 minutes of bad solar data until just now.

Apr 16

21:00 UT The glitches from yesterday re-appeared, so Natsuha stopped the observation and rebooted the correlator at ~21:00 UT. No good scans could be obtained, so she gave up and resumed the solar observation at 23:28 UT. Another attempt will be made at 3 UT next day when 3c273 will rise.

Apr 17

14:50 UT The 3c273 scans for the delay reset last night were unsuccessful due to continuous high wind. This morning Natsuha tried again with 2136+006. The wind was still high but a couple of scans went through without windscram. However, the delay values were very difficult to find, perhaps still due to mild wind speed (9~13 mph). A set of values (mostly zeros) from 131248 scan were sent in at ~14:30 UT.

Another thing - the gliches for which this correlator reboot was done in the first place reappeared this morning. This indicates that the correlator reboot no longer can solve this problem. The delay reset will still be attempted at the earliest possible moment since the glitches appear on only selected bands and other bands' data are fine.

Apr 23

21:26 UT Although not recorded here, we discovered last week that the Rubidium clock was not working (the 10 MHz output was only 138 Hz!), so we have not been observing up to now, and the couple of days prior to Apr. 17 were no good also. We made an emergency order for a new clock, and just now Kjell replaced the old with the new. Right now I have rebooted the correlator and am tracking 3C84. A new scan will start at 21:32 UT, which I can use to update the delays and we should be good to go.